Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?
2x-4=20 is what I have as the start of this equation, is this correct so far?
I appreciate your time on this.
Thank You,
Barb Neely

Answer by skigal545(2)

first you must come up with what x and y represent.
x= his sister
y= Richardo
Your first equation, 2x-4=20 is NOT correct because this is saying that 2 times the age minus four years is equal to both of their ages together. What it should say is 2x-4=y meaning two times his sisters age minus four years is equal to Richardo's age. The second equation would be their ages added together equals 20. so the equation would be x+y=20. From here you would get x or y alone in the second problem and plug it into the first.
x+y=20
x-x+y=20-x
y=20-x and plug this in.